OpenStudy (anonymous):

This is Algebra 2... (-d) to the fourth given that a=3, b=2a-4, c=a+b, and d= -(a+2b)

Directrix (directrix):

Is that (-d)^4 or - (d)^4 ?

OpenStudy (anonymous):

d=-a-2b d=-3-2(2(3)-4) d= -3-12+8 d=-7 -7^4=+2401 .. sorry if i made any minor mistakes but should be right